A sample of hydrogen (H2) gas was collected over water at 60°C.

If the total volume of gas collected was 60.5 mL and the atmospheric pressure was 1.1 bar, what mass (in g) of hydrogen gas was collected above the water?

The vapour pressure of water at 60°C is 150 torr.

Answer:



1
Expert's answer
2022-02-08T11:12:02-0500

"PV= nRT"

Pressure= 1.1-0.2=0.9

Volume= 0.0605L

T= 333.15k

0.9× 0.0605= n× 8.314× 333.15

n = 0.05445/2769.8= 1.97× 10^-5

Mass= 1.97× 10^-5g